In a normal distribution 26% of the scores are below 40 and 7% are above 80. Find out the mean and standard deviation of the distribution?
Answer by stanbon(75887)

In a normal distribution 26% of the scores are below 40 and 7% are above 80. Find out the mean and standard deviation of the distribution?
------
Find the z-value with a left tail of 0.26
and the z-value with a left tail of 0.93.
---
invNorm(0.26) = -0.6433
invNorm(0.93) = 1.4758
----
Create two equations using x = z*s + u
40 = -0.6433*s + u
80 = 1.4758*s + u
----
Subtract and solve for "s":
40 = 2.1246s
s = 40/2.1246 = 18.8272
-------
Solve for "u":
u = 40 + 0.6433*18.8272 = 52.11
==================================
Cheers,
Stan H.
=============